OpenStudy (anonymous):

f(x) = -X^2 + 6x -11 solve for f(t-3)

OpenStudy (luigi0210):

Plug in t-3 so: \[\LARGE f(t-3)=-(t-3)^2+6(t-3)-11\] Now just distribute, combine like terms, etc.

OpenStudy (anonymous):

I got -t^2 +6t -38 but the solution states its -t^2 +12t -38, where did the 12t come from?

OpenStudy (anonymous):

or is the solution on the pdf incorrect?

OpenStudy (luigi0210):

Show me your steps and how you approached the problem, then I'll let you know if you did something wrong.

OpenStudy (anonymous):

-(t-3)^2 = -(t^2 + 9) = -t^2 - 9 6(t -3) = 6t -18 -9 + -18 + -11 = -38 f(t-3) = -t^2 +6t -38

OpenStudy (luigi0210):

Ah, there's your mistake: \[\LARGE (a+b)^2 \ne a^2+b^2\]

OpenStudy (luigi0210):

or if it was minus: \[\LARGE (a \pm b)^2 \ne a^2 \pm b^2\]

OpenStudy (luigi0210):

When doing \[\LARGE (t-3)^2\] \[\LARGE (t-3)^2 \ne (t^2)+(-3)^2\] It's actually: \[\LARGE (t-3)^2=(t-3)(t-3)\] So distribute this and get: \[\LARGE t^2-6t+9\]

OpenStudy (luigi0210):

Now just distribute that negative and finish it: \[\LARGE -(t^2-6t+9)+6(t-3)-11\]

OpenStudy (anonymous):

-t^2 +6t -9 +6t -18 -11 = -t^2 +12t -38

OpenStudy (luigi0210):

There ya go :)

OpenStudy (anonymous):

thank you how do you write the formula so neat when you reply?

OpenStudy (luigi0210):

At the bottom of this reply box there are four buttons right? Use the one that has the sigma and "Equation"

OpenStudy (luigi0210):

But I've been practicing it so much that I can just write it out in plain text without it.

OpenStudy (anonymous):

Sorry for bothering with simple calculus haven't been in school for 7 years and am going back to school so am brushing up.

OpenStudy (luigi0210):

Oh, it's no problem at all, always happy to help~

OpenStudy (luigi0210):

Also: \[\LARGE \mathbb{\color{green}{WELCOME~~TO~~OPENSTUDY~\ddot\smile}}\]

OpenStudy (luigi0210):

If you have any questions on the site feel free to shoot me a message, hope you enjoy your time here :)

OpenStudy (anonymous):

Am going to close this question but could you break down (a+b) 2 ≠a 2 +b 2 a bit more?

OpenStudy (luigi0210):

Oh, and if you haven't already, it would be helpful to read the CoC to avoid any trouble: http://openstudy.com/code-of-conduct

OpenStudy (luigi0210):

And break it down? As in explain it?

OpenStudy (anonymous):

(t−3)^2 =(t−3)(t−3) yes I do understand where t^2 and 9 come from but where is 6t?

OpenStudy (luigi0210):

Well when you have \[\LARGE (t-3)^2\] It's saying t-3 squared, as in multiply it by itself. so \[\LARGE (t-3)^2=(t-3)(t-3)\] And in that cause you would use the FOIL properties: F-First Outer Inner Last